Agustin Industries is a division of a major corporation. Data concerning the most recent year appears below. Sales Net operatThe following data are for the Akron Division of Consolidated Rubber, Inc.: Sales Net operating income Average operating asseMultiple Choice Ο 30% Ο 12% Ο 6% Ο 15%
Answer #1

Return on investment = Net operating income /Average operating assets

= 1,581,000/5,000,000

= 31.62%

Option A

Residual income = Net operating income - Required return

15,000 = 45,000 - (x * 250,000)

(x * 250,000) = 30,000

x = 30,000/250,000 = 12%

Option B